# MBSS-T1: Model-Based Self-Supervised Motion Correction for Robust Cardiac T1 Mapping

### v1

@@ 3.2 Loss Functions
$\begin{split}\mathcal{L}_{fit}(T_{1},M_{0},t_{i=0}^{N-1},\Phi)=&\sum_{i=0}^{N-1}(S_{i}-R_{i})^{2}\\ =&\sum_{i=0}^{N-1}\left(M_{0}\cdot(1-2\cdot e^{\frac{-t_{i}}{T_{1}}})-R_{i}\right)^{2}\end{split}$ (8)
@@
$\mathcal{L}_{smooth}(\Phi)=\sum_{i=0}^{N-1}\frac{1}{\Omega}\sum_{p\in\Omega}\left\lVert\nabla\phi_{i}(p)\right\rVert^{2}$ (9)
Here, $\Omega$ denotes the domain of the velocity field, and $p$ represents the voxel locations within this domain. By minimizing this term, the model promotes smoother deformation fields, reducing unrealistic distortions.
The last term ($\mathcal{L}_{Seg}$), inspired by [2], encourages anatomically-correct deformation fields by maximizing the Dice score between the deformed myocardium segmentations and by maximizing the Dice score between the deformed left ventricle segmentations. We denote $\mathcal{L}_{Seg}$ as follows:
$\mathcal{L}_{Seg}=\mathcal{L}_{MyoSeg}+\mathcal{L}_{LVSeg}$ (10)
In our approach, we focus on the segmentations that are retained in the set $G$, which contains only those segmentations with high confidence. The $\mathcal{L}_{MyoSeg}$ and $\mathcal{L}_{LVSeg}$ segmentation loss functions are therefore defined using only these high-confidence segmentations:
@@
where $MyoSeg_{i}$ is the $i$-th binary segmentation mask of the myocardium, $LVSeg_{i}$ is the $i$-th binary segmentation mask of the left ventricle, and $Seg_{r}$ is the binary segmentation mask of the $r$-th image, which is the reference image. Only the segmentations indexed by $G$, which meet the high-confidence criteria, are used in the calculation of the loss.
@@ 3.3 Implementation details
We implemented our models in PyTorch. Experimentally, we zeroed the $r$-th deformation field and predicted deformation fields only for the rest of the time points. We optimized our hyperparameters using a grid search. The final settings for the loss function parameters were: $\lambda_{1}=1$, $\lambda_{2}=5000$, $\lambda_{3}=80000$, $\gamma=0.99$, and $\alpha=0.9$. We used a batch size of 1 and the ADAM optimizer with a learning rate of $2\cdot 10^{-3}$. Our network overfits to each case. For each case, we trained the network from scratch for a fixed number of 40 iterations, which took approximately 6 seconds. We performed hyperparameter optimization for the baseline methods using a grid search. All experiments were run on an NVIDIA Tesla V100 GPU with 32GB of RAM.
@@ 4.1.1 Free Breathing - STONE Dataset
A publicly available myocardial T1 mapping dataset [4, 5] consisting of 210 subjects, including 134 males and 76 females, aged $57\pm 14$ years. All subjects were diagnosed with or suspected of having cardiovascular diseases. The imaging was performed on a 1.5T MRI scanner (Philips Achieva) equipped with a 32-channel cardiac coil, using an ECG-triggered, free-breathing, slice-interleaved T1 mapping sequence (STONE) [29]. The acquisition parameters were as follows: field of view (FOV) $=360\times 351$ [mm2] and voxel size $=2.1\times 2.1\times 8$ [mm3]. For each subject, 5 slices were captured from the base to the apex in the short-axis view at 11 distinct time points. Additionally, the dataset included manual expert segmentations of the myocardium [4]. The images were resized to $160\times 160$ pixels per time point, and min-max normalization was applied to the entire sequence to standardize the image intensities.
@@ 4.1.2 Free Breathing and Breath Hold - MOLLI Dataset
Data was acquired on a 3T Siemens MRI system (PRISMA, Siemens Healthineers, Erlangen, Germany) equipped with a 32-channel body coil, using an ECG-triggered, free-breathing, slice-interleaved T1 mapping MOLLI sequence (MyoMaps, SIEMENS Healthcare). The acquisition parameters were as follows: field of view (FOV) $=306\times 360$ [mm2], flip angle of 35°, and an 8 mm gap between slices. For each subject, 5 slices were captured from the base to the apex in the short-axis view at 8 distinct time points. The dataset consists of cardiac images obtained from 19 patients, with each patient undergoing two scans: one during breath-hold and another during free breathing. Additionally, for five of the patients, extra breath-hold and free-breathing scans were performed for test-retest analysis.
@@ 4.2 Experiments
Experimental setup: Our MBSS-T1 method was compared against two variations: a multi-image registration model using a mutual-information-based loss function (REG-MI) [12], and PCMC-T1, which incorporates physical constraints into the motion correction process within an unsupervised framework [13]. Additionally, two state-of-the-art deep learning algorithms for medical image registration were included in the comparison: pairwise probabilistic diffeomorphic VoxelMorph with a mutual-information-based loss [3], and pairwise SynthMorph [14]. For the self-supervised approaches, MBSS-T1 and SynthMorph, the entire STONE dataset was utilized for inference since no training phase was required. For the unsupervised approaches (PCMC-T1, REG-MI, VoxelMorph) and the nnUNet pretrained segmentation network (used for MBSS-T1), a 5-fold cross-validation strategy was implemented. In each fold, the 210 subjects were divided into 80% for training and 20% for testing. T1 maps generated directly from the acquired images were also used for comparison by applying a standard T1 mapping procedure. In this process, a T1 relaxation curve was fitted to the MRI data using the conventional inversion recovery model, without any pre-motion correction techniques. Fitting was performed on a per-pixel basis across the series of T1-weighted images, resulting in maps that represent the T1 relaxation times. These maps referred to as ’w.o motion correction’, serve as a baseline to evaluate the impact of motion correction on the accuracy and quality of the T1 mapping. The T1 maps generated by our MBSS-T1 model were quantitatively evaluated against those produced both before (’w.o motion correction’) and after deep-learning-based image registration as a pre-processing step. The evaluation metrics used were the $R^{2}$ of the model fit to the observed data in the myocardium, the Dice score, and the Hausdorff distance values of the myocardium segmentations.
Clinical impact: The clinical impact of our method was assessed through a semi-quantitative ranking of the T1 maps for motion artifacts. This assessment was conducted by an expert cardiac MRI radiologist with 10 years of experience (I.V), who was blinded to the methods used to generate the maps. A random selection of 29 cases (each with 5 slices) from the test set, along with their corresponding T1 maps, was evaluated. The radiologist rated each slice, assigning a score of 1 to slices without visible motion artifacts and a score of 0 to those with artifacts. Overall patient scores were calculated by summing the scores of the slices. The highest possible score per subject was 5, indicating no motion artifacts in all slices, while the lowest was 0, indicating artifacts in all slices. Statistical significance was evaluated using the repeated measures ANOVA test, with p$<$0.05 considered significant.
@@
Clinical Impact: As in Exp 1, the clinical impact of our method was assessed through a semi-quantitative ranking of the T1 maps for the presence of motion artifacts. This assessment was conducted on both free-breathing and breath-hold scans to determine how effectively each method mitigated motion artifacts. The primary evaluation metric was a clinical score, assigned by an experienced cardiac MRI radiologist with 10 years of experience, based on the clarity and quality of the T1 maps produced by each method.
@@
Experimental Setup and Objective: The objective of this experiment was to assess the consistency and reproducibility of T1 mapping when motion correction is applied. This analysis aimed to evaluate whether motion correction techniques improve the reliability of myocardial T1 measurements across repeated scans, particularly in the presence of motion. The subset included data from five patients, each undergoing two free-breathing scans and two breath-hold scans. The goal was to compare the consistency of T1 measurements across these scans using different motion correction methods.
Quantitative Evaluation: To quantitatively assess the reproducibility of T1 measurements, we calculated the intraclass correlation coefficient (ICC3) [19] for mean myocardial T1 values across three different methods: without motion correction, Siemens MyoMaps, and MBSS-T1. The analysis was conducted using data from five patients in the MOLLI dataset, each of whom underwent two free-breathing scans and two breath-hold scans. For each scan and method, mean myocardial T1 values were computed for four regions of interest (ROIs) – upper, right, lower, and left – on the mid-slice of the T1 map. The ICC3 metric was then calculated for all paired test-retest ROIs to assess the agreement between repeated measurements. A high ICC3 value indicates strong agreement between the test and retest measurements, suggesting robustness and reproducibility in quantifying myocardial T1 values. This metric provides valuable insights into the reliability and consistency of T1 measurements across different methods and scanning conditions. By quantifying the agreement between test and retest measurements across different ROIs and scanning conditions, a comprehensive understanding of each method’s performance in clinical practice can be achieved. This analysis also enables the evaluation of the effectiveness of motion correction techniques in improving measurement accuracy and reproducibility.
@@ 5 Results
Quantitative evaluation: Table 1 summarizes our results for the test sets across all folds, encompassing a total of 210 patients. The MBSS-T1 method outperformed other state-of-the-art registration methods in terms of $R^{2}$, Dice score, and Hausdorff distance. Specifically, MBSS-T1 achieved an $R^{2}$ value of $0.975\pm 0.05$, which represents an improvement of approximately 3% compared to the next best method, PCMC-T1, with an $R^{2}$ of $0.955\pm 0.078$. Additionally, MBSS-T1 produced a Dice score of $0.89\pm 0.075$, surpassing SynthMorph’s $0.88\pm 0.149$, and reduced the Hausdorff distance to $6.43\pm 5.54$ mm, which is a 31% improvement compared to SynthMorph’s $8.59\pm 9.98$ mm. These results suggest that our approach enhances the physical plausibility of deformations through signal relaxation and anatomical consistency. Our results are particularly noteworthy considering that we made comparisons with methods requiring a full training process.
Clinical impact: Figure 3 presents several representative cases. The maps generated by the MBSS-T1 method demonstrate the highest quality, as evidenced by superior metrics. The rightmost column of Table 1 summarizes the clinical impact assessment results for the MBSS-T1 method, which achieved the highest quality score of $4.33\pm 0.54$. This score reflects an improvement of approximately 10% over the next best method, PCMC-T1, which had a score of $3.93\pm 0.78$. The difference in radiologist grading was statistically significant (p $\ll 10^{-5}$). Furthermore, as demonstrated in Fig. 5, the MBSS-T1 method achieves a closer fit to the expected magnetization curve (with $R^{2}=0.99$) and reduces artifacts, which may contribute to improved diagnostic confidence. These results suggest the MBSS-T1 method’s potential utility in clinical settings where motion artifacts can impact the accuracy of T1 mapping.
@@
The images in Fig. 4 illustrate the reduction in motion-related distortions achieved by the MBSS-T1 method. This improved motion correction results in clearer and more reliable T1 maps with fewer artifacts.
@@
For both mean and median myocardial T1 values, the MBSS-T1 method consistently exhibited higher ICC3 values compared to the other methods. Specifically, for the mean T1 values, the MBSS-T1 method attained ICC3 values of $0.23$ for free-breathing (FB), $0.42$ for breath-hold (BH), and $0.31$ for combined FB and BH scans. In contrast, the w.o motion correction and Siemens Myomap methods displayed significantly lower ICC3 values, suggesting less agreement between test and retest measurements. Similarly, for the median T1 values, the MBSS-T1 method achieved ICC3 values of $0.26$ (FB), $0.38$ (BH), and $0.3$ (FB+BH), surpassing the other methods.
These findings suggest that the MBSS-T1 method offers more reliable and consistent myocardial T1 measurements compared to the other methods. However, it is important to note that while the MBSS-T1 method shows the highest ICC3 values, the scores indicate only moderate reliability. This means that although the MBSS-T1 method improves the robustness and reproducibility of T1 mapping under both free-breathing and breath-hold conditions, further improvements are needed to achieve high reliability.
@@
Figure 4: Comparison of T1 mapping results for four patients (rows) using three motion correction methods: Original (left), Siemens MyoMap (middle), and MBSS-T1 (right). MBSS-T1 provides superior motion correction, resulting in clearer T1 maps with fewer artifacts.
@@ 6 Conclusions
This work presents a self-supervised method for motion-robust cardiac T1 mapping. The method utilizes a self-supervised learning approach combined with a pre-trained segmentation framework, enhanced by incorporating a signal decay model within the network’s architecture along with an anatomical constraint. This approach enables the generation of corrected T1 maps across various imaging protocols without requiring protocol-specific data collection or extensive training. The main contribution of this work lies in the novel combination of signal decay and anatomical constraints within the registration process, providing a powerful tool for improving motion correction and expanding the reach of quantitative MRI techniques in clinical practice.
The method demonstrated quantitative improvements in fit quality, Dice score, and Hausdorff distance compared to baseline methods. A qualitative evaluation by an expert cardiac radiologist indicated the clinical relevance of the generated T1 maps, showing potential for accurate T1 estimation in patients unable to perform breath-hold during MRI scans. This approach may extend the applicability of quantitative cardiac T1 mapping across a broader patient population, ensuring reliable imaging even under challenging conditions.
In addition to its effectiveness in T1 mapping, the proposed approach holds significant potential for application to other quantitative MRI problems. By integrating signal decay modeling with anatomical constraints, this method offers a robust framework that could be adapted to correct motion and improve accuracy in various quantitative MRI sequences, such as T2 mapping, diffusion MRI, and beyond. Furthermore, the framework is flexible and could be further enhanced by replacing the backbone with more advanced architectures, such as transformer-based models. This substitution could leverage the global attention mechanisms inherent in transformers, potentially improving the capture of complex anatomical structures and further enhancing the accuracy of the generated maps.

### v2

@@ 3.2 Loss Functions
$\begin{split}\mathcal{L}_{fit}(T_{1},M_{0},t_{i=0}^{N-1},\Phi)=&\sum_{i=0}^{N-1}(S_{i}-R_{i})^{2}\\ =&\sum_{i=0}^{N-1}\left(M_{0}\cdot(1-2\cdot e^{\frac{-t_{i}}{T_{1}}})-R_{i}\right)^{2}\,.\end{split}$ (8)
@@
$\mathcal{L}_{smooth}(\Phi)=\sum_{i=0}^{N-1}\frac{1}{\Omega}\sum_{p\in\Omega}\left\lVert\nabla\phi_{i}(p)\right\rVert^{2}\,.$ (9)
Here, $\Omega$ denotes the domain of the velocity field, and $p$ represents the voxel locations. By minimizing this term, the model promotes smoother deformation fields, reducing unrealistic distortions.
The last term ($\mathcal{L}_{Seg}$), inspired by [2], encourages anatomically correct deformation fields by maximizing the Dice score between the deformed myocardium segmentations and by maximizing the Dice score between the deformed left ventricle segmentations. We denote $\mathcal{L}_{Seg}$ as follows:
$\mathcal{L}_{Seg}=\mathcal{L}_{MyoSeg}+\mathcal{L}_{LVSeg}\,.$ (10)
Our approach focuses on the segmentations retained in the set $G$, which contains only those segmentations with high confidence. The $\mathcal{L}_{MyoSeg}$ and $\mathcal{L}_{LVSeg}$ segmentation loss functions are therefore defined using only these high-confidence segmentations:
@@
where $MyoSeg_{i}$ is the $i$-th binary segmentation mask of the myocardium, $LVSeg_{i}$ is the $i$-th binary segmentation mask of the left ventricle, and $Seg_{r}$ is the binary segmentation mask of the $r$-th image, which is the reference image. Only the segmentations indexed by $G$, which meet the high-confidence criteria, are used to calculate the loss.
@@ 3.3 Implementation details
We implemented our models in PyTorch. Experimentally, we zeroed the $r$-th deformation field and predicted deformation fields only for the rest of the time points. We optimized our hyperparameters using a grid search. The final settings for the loss function parameters were: $\lambda_{1}=1$, $\lambda_{2}=5000$, $\lambda_{3}=80000$, $\gamma=0.99$, and $\alpha=0.9$. We used a batch size of 1 and the ADAM optimizer with a learning rate of $2\cdot 10^{-3}$. Our network overfits each case. For each case, we trained the network from scratch for a fixed number of 40 iterations, which took approximately 6 seconds. We performed hyperparameter optimization for the baseline methods using a grid search. All experiments were run on an NVIDIA Tesla V100 GPU with 32GB of RAM.
@@ 4.1.1 Free Breathing - STONE Dataset
A publicly available myocardial T1 mapping dataset [4, 5] consists of 210 subjects, 134 males and 76 females, aged $57\pm 14$ years. All subjects were diagnosed with or suspected of having cardiovascular diseases. The imaging was performed on a 1.5T MRI scanner (Philips Achieva) equipped with a 32-channel cardiac coil, using an ECG-triggered, free-breathing, slice-interleaved T1 mapping sequence (STONE) [30]. The acquisition parameters were as follows: field of view (FOV) $=360\times 351$ [mm2] and voxel size $=2.1\times 2.1\times 8$ [mm3]. Five slices were captured from the base to the apex in the short-axis view for each subject at 11 distinct time points. Additionally, the dataset included manual expert segmentations of the myocardium [4]. The images were resized to $160\times 160$ pixels per time point, and min-max normalization was applied to the entire sequence to standardize the image intensities.
@@ 4.1.2 Free Breathing and Breath Hold - MOLLI Dataset
Data was acquired on a 3T Siemens MRI system (PRISMA, Siemens Healthineers, Erlangen, Germany) equipped with a 32-channel body coil, using an ECG-triggered, free-breathing, slice-interleaved T1 mapping MOLLI sequence (MyoMaps, SIEMENS Healthcare). The acquisition parameters were as follows: field of view (FOV) $=306\times 360$ [mm2], flip angle of 35°, and an 8 mm gap between slices. Five slices were captured from the base to the apex in the short-axis view for each subject at eight distinct time points. The dataset consists of cardiac images obtained from 19 patients undergoing two scans: one during breath-hold and another during free breathing. Additionally, extra breath-hold and free-breathing scans were performed for five patients for test-retest analysis.
@@ 4.2 Experiments
Experimental setup: Our MBSS-T1 method was compared against two variations: a multi-image registration model using a mutual-information-based loss function (REG-MI) [12], and PCMC-T1, which incorporates physical constraints into the motion correction process within an unsupervised framework [13]. Additionally, two state-of-the-art deep learning algorithms for medical image registration were included in the comparison: pairwise probabilistic diffeomorphic VoxelMorph with a mutual-information-based loss [3] and pairwise SynthMorph [15]. For the self-supervised approaches, MBSS-T1 and SynthMorph, the entire STONE dataset was utilized for inference since no training phase was required. For the unsupervised approaches (PCMC-T1, REG-MI, VoxelMorph) and the nnUNet pretrained segmentation network (used for MBSS-T1), a 5-fold cross-validation strategy was implemented. In each fold, the 210 subjects were divided into 80% for training and 20% for testing. T1 maps generated directly from the acquired images were also used for comparison by applying a standard T1 mapping procedure. A T1 relaxation curve was fitted to the MRI data using the conventional inversion recovery model without any pre-motion correction techniques. Fitting was performed per pixel across the series of T1-weighted images, resulting in maps representing the T1 relaxation times. These maps, referred to as ’w.o motion correction,’ serve as a baseline to evaluate the impact of motion correction on the accuracy and quality of the T1 mapping. The T1 maps generated by our MBSS-T1 model were quantitatively evaluated against those produced both before (’w.o motion correction’) and after deep-learning-based image registration as a pre-processing step. The evaluation metrics used were the $R^{2}$ of the model fit to the observed data in the myocardium, the Dice score, and the Hausdorff distance values of the myocardium segmentations.
Clinical impact: The clinical implications of our method were assessed through a semi-quantitative ranking of the T1 maps for motion artifacts. This assessment was conducted by an expert cardiac radiologist with five years of experience (I.V), blinded to the methods used to generate the maps. A random selection of 29 cases (each with five slices) from the test set and their corresponding T1 maps was evaluated. The radiologist rated each slice, assigning a score of 1 to slices without significant visible motion artifacts and a score of 0 to those with artifacts. Overall patient scores were calculated by summing the scores of the slices. The highest possible score per subject was 5, indicating no motion artifacts in all slices, while the lowest was 0, indicating artifacts in all slices. Statistical significance was evaluated using the repeated measures ANOVA test, with p$<$0.05 considered significant.
@@
Clinical Impact: As in Exp 1, the clinical impact of our method was assessed through a semi-quantitative ranking of the T1 maps for the presence of motion artifacts. This assessment was conducted on free-breathing and breath-hold scans to determine how effectively each method mitigated motion artifacts. The primary evaluation metric was a clinical score assigned by an experienced cardiac radiologist with five years of experience based on the clarity and quality of the T1 maps produced by each method.
@@
Experimental Setup and Objective: This experiment aimed to assess the consistency and reproducibility of T1 mapping when motion correction is applied. This analysis aimed to evaluate whether motion correction techniques improve the reliability of myocardial T1 measurements across repeated scans, particularly in the presence of motion. The subset included data from five patients undergoing two free-breathing and two breath-hold scans. The goal was to compare the consistency of T1 measurements across these scans using different motion correction methods.
Quantitative Evaluation: To quantitatively assess the reproducibility of T1 measurements, we calculated the intraclass correlation coefficient (ICC3) [20] for mean myocardial T1 values across three different methods: without motion correction, Siemens MyoMaps, and MBSS-T1. The analysis was conducted using data from five patients in the MOLLI dataset, each of whom underwent two free-breathing and two breath-hold scans. For each scan and method, mean myocardial T1 values were computed for four regions of interest (ROIs) – upper, right, lower, and left – on the mid-slice of the T1 map. The ICC3 metric was then calculated for all paired test-retest ROIs to assess the agreement between repeated measurements. A high ICC3 value indicates strong agreement between the test and retest measurements, suggesting robustness and reproducibility in quantifying myocardial T1 values. This metric provides valuable insights into the reliability and consistency of T1 measurements across different methods and scanning conditions. A comprehensive understanding of each method’s performance in clinical practice can be achieved by quantifying the agreement between test and retest measurements across different ROIs and scanning conditions. This analysis also enables the evaluation of the effectiveness of motion correction techniques in improving measurement accuracy and reproducibility.
@@ 5 Results
Quantitative evaluation: Table 1 summarizes our results for the test sets across all folds, encompassing a total of 210 patients. The MBSS-T1 method outperformed other state-of-the-art registration methods in terms of $R^{2}$, Dice score, and Hausdorff distance. Specifically, MBSS-T1 achieved an $R^{2}$ value of $0.975\pm 0.05$, which represents an improvement of approximately 3% compared to the next best method, PCMC-T1, with an $R^{2}$ of $0.955\pm 0.078$. Additionally, MBSS-T1 produced a Dice score of $0.89\pm 0.075$, surpassing SynthMorph’s $0.88\pm 0.149$, and reduced the Hausdorff distance to $6.43\pm 5.54$ mm, which is a 31% improvement compared to SynthMorph’s $8.59\pm 9.98$ mm. These results suggest that our approach enhances the physical plausibility of deformations through signal relaxation and anatomical consistency. Our results are particularly noteworthy, considering we compared methods requiring a full training process.
Clinical impact: Figure 3 presents several representative cases. The maps generated by the MBSS-T1 method demonstrate the highest quality, as evidenced by superior metrics. The rightmost column of Table 1 summarizes the clinical impact assessment results for the MBSS-T1 method, which achieved the highest quality score of $4.33\pm 0.54$. This score reflects an improvement of approximately 10% over the following best method, PCMC-T1, which had a score of $3.93\pm 0.78$. The difference in radiologist grading was statistically significant (p $\ll 10^{-5}$). Furthermore, as demonstrated in Fig. 5, the MBSS-T1 method achieves a closer fit to the expected magnetization curve (with $R^{2}=0.99$) and reduces artifacts, which may contribute to improved diagnostic confidence. These results suggest the MBSS-T1 method’s potential utility in clinical settings where motion artifacts can impact the accuracy of T1 mapping.
@@
The images in Fig. 4 illustrate the reduction in motion-related distortions achieved by the MBSS-T1 method. This improved motion correction results in more precise and reliable T1 maps with fewer artifacts.
@@
For both mean and median myocardial T1 values, the MBSS-T1 method consistently exhibited higher ICC3 values than the other methods. Specifically, for the mean T1 values, the MBSS-T1 method attained ICC3 values of $0.23$ for free-breathing (FB), $0.42$ for breath-hold (BH), and $0.31$ for combined FB and BH scans. In contrast, the w.o motion correction and Siemens Myomap methods displayed significantly lower ICC3 values, suggesting less agreement between test and retest measurements. Similarly, for the median T1 values, the MBSS-T1 method achieved ICC3 values of $0.26$ (FB), $0.38$ (BH), and $0.3$ (FB+BH), surpassing the other methods.
These findings suggest that the MBSS-T1 method offers more reliable and consistent myocardial T1 measurements than the other methods. However, it is important to note that while the MBSS-T1 method shows the highest ICC3 values, the scores indicate only moderate reliability. Although the MBSS-T1 method improves the robustness and reproducibility of T1 mapping under free-breathing and breath-hold conditions, further improvements are needed to achieve high reliability.
@@
Figure 4: Comparison of T1 mapping results for four patients (rows) using three motion correction methods: Original (left), Siemens MyoMap (middle), and MBSS-T1 (right). MBSS-T1 provides superior motion correction, resulting in clearer T1 maps with fewer artifacts and better differentiation between the blood and myocardium.
@@ 6 Conclusions
This work presents a self-supervised method for motion-robust cardiac T1 mapping. The method utilizes a self-supervised learning approach combined with a pre-trained segmentation framework, enhanced by incorporating a signal decay model within the network’s architecture and an anatomical constraint. This approach enables the generation of corrected T1 maps across various imaging protocols without requiring protocol-specific data collection or extensive training. The main contribution of this work lies in the novel combination of signal decay and anatomical constraints within the registration process, providing a powerful tool for improving motion correction and expanding the reach of quantitative MRI techniques in clinical practice.
Compared to baseline methods, the method demonstrated quantitative improvements in fit quality, Dice score, and Hausdorff distance. A qualitative evaluation by an expert cardiac radiologist indicated the clinical relevance of the generated T1 maps, showing potential for accurate T1 estimation in patients unable to perform breath-hold during MRI scans. This approach may extend the applicability of quantitative cardiac T1 mapping across a broader patient population, ensuring reliable imaging even under challenging conditions.
In addition to its effectiveness in T1 mapping, the proposed approach holds significant potential for application to other quantitative MRI problems. By integrating signal decay modeling with anatomical constraints, this method offers a robust framework that could be adapted to correct motion and improve accuracy in various quantitative MRI sequences, such as T2 mapping, diffusion MRI, and beyond. Furthermore, the flexible framework could be further enhanced by replacing the backbone with more advanced architectures, such as transformer-based models. This substitution could leverage the global attention mechanisms inherent in transformers, potentially improving the capture of complex anatomical structures and further enhancing the accuracy of the generated maps.